Find the area of the surface generated by revolving the curve


y

=

5

x

,

y=5x,

       2

x

4

2≤x≤4



around the x

x

-axis.


1
Expert's answer
2021-03-10T11:31:44-0500

"y\ =\ 5\sqrt x\"

"\frac{dy}{dx}=\frac{5}{2\sqrt x}"

"\left(\frac{dy}{dx}\right)^2=\frac{25}{4x}"

"\int_{2}^{4}{2\pi y\ ds}=\ \int_{2}^{4}{2\pi5\sqrt x\sqrt{1+\ \frac{25}{4x}}dx}"

"10\pi\int_{2}^{4}{\sqrt{x+\ \frac{25}{4}}dx}"

"10\pi\ \left(x+\frac{25}{4}\right)^\frac{3}{2}\ast\frac{2}{3}"

"\frac{20\pi}{3}\left\{\left(\frac{41}{4}\right)^\frac{3}{2}-\left(\frac{33}{4}\right)^\frac{3}{2}\right\}"

"60.79794\pi"